OpenStudy (sarahezzmck):

Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5, will someone help me with this please?

OpenStudy (mathstudent55):

Is this the equation? \( \dfrac{2}{x + 2} + \dfrac{1}{5} = \dfrac{6}{x + 5} \)

OpenStudy (sarahezzmck):

Yes, it is :)

OpenStudy (sarahezzmck):

I was thinking that it may be x=13 because 13+2 is 15 and in order to get 15 you have to multiply 3 by 5. And to get 6 you have to multiply 6 by 3. @mathstudent55

OpenStudy (sarahezzmck):

Does that make any sense???